# Complete Security and Privacy for AI Inference in Decentralized Systems

### v1

@@ Abstract
The need for data security and model integrity has been accentuated by the rapid adoption of AI and ML in data-driven domains including healthcare, finance, and security. Large models are crucial for tasks like diagnosing diseases and forecasting finances but tend to be delicate and not very scalable. Decentralized systems solve this issue by distributing the workload and reducing central points of failure. Yet, data and processes spread across different nodes can be at risk of unauthorized access, especially when they involve sensitive information. Nesa solves these challenges with a comprehensive framework using multiple techniques to protect data and model outputs. This includes zero-knowledge proofs for secure model verification. The framework also introduces consensus-based verification checks for consistent outputs across nodes and confirms model integrity. Split Learning divides models into segments processed by different nodes for data privacy by preventing full data access at any single point. For hardware-based security, trusted execution environments are used to protect data and computations within secure zones. Nesa’s state-of-the-art proofs and principles demonstrate the framework’s effectiveness, making it a promising approach for securely democratizing artificial intelligence.
@@ 1 Introduction
Artificial Intelligence (AI), particularly machine learning (ML), has made significant strides in recent decades. Since the advent of large language models (LLMs) such as ChatGPT [1], Claude [2], Gemini [3], LLaMA [4] and diffusion models [5] such as DALLE-3 [6] and Sora [7], foundation models have garnered considerable attention. While these foundation models exhibit intriguing properties such as in-context learning and chain-of-thought reasoning, concerns about their security and privacy have emerged, especially in distributed or decentralized computing scenarios. For instance, in ML as a service (MLaaS), ensuring the integrity of inference results is paramount. Service providers must demonstrate to customers that the output stems from inputting the customer’s prompt into a verified large language model, like GPT-4, and generating the response through model execution, rather than relying on human writers or less advanced models, such as GPT-3.5. This requirement is referred to as model security or model integrity. Meanwhile [8], service providers are reluctant to release their model weights, preferring to keep them confidential. In distributed or decentralized inference applications, a central server divides a foundation model into distinct segments, each managed by a different party. When an inference query arises, each party executes its segment independently. It is key to ascertain the trustworthiness of each party’s execution before aggregating the outputs.
Meanwhile, data privacy issues also arise in decentralized inference [9], where users need to ensure that their input data are not directly visible to the decentralized nodes that carry the AI inference for them. This can be extremely important in “critical inference” settings where sensitive information, such as medical records, financial data, or security information, is processed. In healthcare, where AI models analyze medical images like MRI or CT scans to diagnose diseases. Patient data is highly sensitive and must be protected. Ensuring data privacy in such scenarios is crucial to protect user data from unauthorized access and potential misuse.
At Nesa, we aim to address security and privacy concerns for AI’s responsible and effective deployment across various domains. In a nutshell, we design hybrid approaches to security and privacy [10] for the above challenges. The selection of different security and privacy methods depends on the specific use cases and the need for varying levels of security and privacy. Based on the different levels of needs, users can flexibly choose the best option for security and privacy. Here, we discuss two key scenarios at Nesa and our security and privacy approaches to them. Also, we discuss how we adapt a hardware-based trusted execution environment (TEE) as an orthogonal approach for security and privacy.
@@
Scenario 1: Critical inference refers to scenarios where the results of AI inference are extremely significant, necessitating the highest levels of security and privacy, even if it means slower speed and higher cost. In these situations, the accuracy and confidentiality of the inference outcomes are important, and users are willing to endure longer processing times to ensure their data is fully protected. Examples include healthcare diagnostics, where AI analyzes medical images to detect diseases, and financial decision-making, where AI evaluates large transactions or investment strategies. In both cases, the sensitive nature of the data demands robust privacy measures, with stakeholders prioritizing data security over rapid results to prevent unauthorized access and ensure the integrity of the process. Under this scenario, Nesa adapts and innovates two leading technologies, including zero-knowledge machine learning (ZKML) [11] for model integrity verification and homomorphic encryption (HE) for data encryption and user privacy protection. Admittedly, the original design of ZKML and HE is computationally heavy, especially for non-linear layers in neural networks such as ReLU layers [12]. To balance effectiveness and efficiency, we apply the state-of-the-art ZKML techniques introduced by [11, 13] to build zero-knowledge Decentralized Proof System (zkDPS) for proof generation and verification processes in foundation models, and we apply HE selectively to critical layers rather than the entire model. Specifically, we propose Sequential Vector Encryption (SVE) based on sequential homomorphic encryption (SHE). This robust encryption scheme obfuscates the output of each operator in a neural network to prevent attackers from extracting sensitive information from intermediate representations. §2 provides more details on our customized zkDPS and SHE solutions for critical inference with the highest protection.
@@
Scenario 2: General inference refers to everyday AI inference tasks with less critical results, allowing for faster processing speeds and lower costs without compromising basic security and privacy standards. In these scenarios, a certain level of protection is still necessary, but the stringent measures required for critical inference are not needed. Examples include routine tasks such as checking the weather, recommending products, or filtering spam emails. Here, the emphasis is on efficiency and speed while maintaining adequate data security to protect user information. Under this scenario, Nesa employs our consensus-based verification (CBV) for security, ensuring the integrity of the inference process, and split learning (SL) [14] for data encryption, which provides a reasonable level of data protection. Notably, we have innovated this verification method to reduce the high redundancy requirements in model verification, and split learning has been used in decentralized AI encryption for the first time. §3 elaborates on our consensus-based verification and split learning solutions for general inference, balancing speed and security effectively.
@@
NTEE: Nesa’s Trusted Execution Environment (TEE). In addition to our innovations around software- and/or algorithm-based security and privacy approaches for the two scenarios, we also design an orthogonal hardware-based approach for both. In a nutshell, TEEs create secure isolation zones within the network’s nodes, protecting user data and private model parameters from unauthorized access [15]. TEEs provide a secure enclave for executing computations, isolated from the rest of the node’s operating environment, ensuring that even if other parts of the node are compromised, the computations within the TEE remain protected. This isolation is critical in decentralized settings where AI models are spread among multiple owners. This hardware-based security measure works as an alternative to algorithm-based approaches, providing a fast and efficient solution for Nesa’s decentralized AI inference tasks. Additionally, we innovate NTEE (Nesa TEEs) optimizing communication among multiple TEEs by establishing direct, secure channels, and implementing heterogeneous TEE scheduling based on the capabilities of each node, whether CPU or GPU-based. NTEE facilitates secure collaboration across multiple nodes, enabling Nesa to maintain high performance, robust security, and data privacy, making it ideal for both critical and general inference scenarios. See details of our TEE implementation and innovation in §4.
@@
Algorithm Critical (§2) zero-knowledge Decentralized Proof System (zkDPS) (§2.1) Sequential Vector Encryption (SVE) (§2.2)
Algorithm General (§3) Consensus-Based Verification (CBV) (§3.1) Split Learning (SL) (§3.2)
Hardware Both TEE model verification (§4) TEE secure channel (§4)
Table 1: Summary of Security and Privacy Solutions (continued)
Choosing Security and Privacy Options. Table LABEL:table:suecurity-all summarizes our innovations around security and privacy. Here, we further discuss the choice of different methods by scenario. Given critical inference scenarios, where the final results have high value and users are willing to wait longer, we suggest using zkDPS for model verification and SHE for data encryption. In contrast, for general inference scenarios, where the results are less critical and faster processing is desired, we recommend using CBV for model verification and SL for data encryption. Additionally, the TEE-based solution, NTEE, can also be leveraged as an orthogonal approach. Once NTEE is employed, it addresses both verification and encryption needs simultaneously. However, it is important to note that GPU TEEs are only available on high-end GPUs, making the trade-off such that TEE solutions are mostly CPU-based. In contrast, algorithm-based solutions can fully leverage GPUs, offering a different balance of performance and security. This approach allows us to balance the trade-offs between security, privacy, and efficiency, ensuring that users can select the most appropriate solution for their specific needs. By leveraging both algorithm-based and hardware-based security measures, Nesa can provide robust and adaptable security and privacy solutions across a wide range of AI applications. This adaptive approach ensures that users can select the most appropriate level of protection for their data and applications, achieving an optimal balance among security, privacy, and performance.
@@ 2 Security and Privacy for Critical Inference
Figure 1: Security flowchart for critical inference, where Nesa designs zkDPS2.1) for model verification and SHE2.2) for data encryption and user privacy protection.
@@
In critical inference scenarios, the significance of AI inference results necessitates the highest levels of security and privacy, as the outcomes directly impact crucial sectors like healthcare and finance. Our strategy for ensuring model verification and integrity is via our Zero-Knowledge Decentralized Proof System (zkDPS), a specialized ZK system for decentralized LLMs that allows one party to prove to another that a statement is true without revealing any information beyond the validity of the statement itself. Notably, it introduces a few new techniques to speed up the ZK process. This approach is detailed in §2.1. For data privacy, we redesign Sequential Homomorphic Encryption (SHE) and propose our Sequential Vector Encryption (SVE), which enables computations on encrypted data without decrypting it, thus ensuring that sensitive data remains protected throughout the inference process. In comparison to classical HE, our approaches can be more efficient in applying implied tensor transformations to selected layers of an AI model. More information on this technique can be found in Section 2.2. Fig. 1 summarizes the security flow of critical inference at Nesa, where the data will be transmitted with SVE for encryption, and the final inference results will be verified by zkDPS for integrity.
2.1 Zero-Knowledge Machine Learning for Model Integrity
2.1.1 Background
@@
2.1.2 Problem Setups
Decentralized Inference. At Nesa, we explore the challenge of decentralized inference for foundation models, which offers numerous benefits. With the advent of 5G technology and improved internet latency, personalized devices such as mobile phones can now participate in crowdsourcing machine learning models. This approach enhances device utilization and eliminates the need for data communication between nodes and a central server, thereby safeguarding data privacy. In the context of Machine Learning as a Service (MLaaS), models remain the private property of individual nodes, allowing them to offer model inference services via APIs without sharing model weights and checkpoints. As foundation models grow in size—such as Meta’s LLaMA-3.1 with 405 billion parameters [4], and future models expected to reach trillion-level sizes—it becomes impractical to load entire models on a single node or server. Additionally, technologies like blockchain are now equipped to support the decentralized inference of foundation models on-chain, further facilitating this approach.
Model Decomposition. A core assumption in decentralized AI inference is that the model used in the inference system can be divided into several parts, each managed by a separate party, or computing node. Each computing node performs its assigned computations and sends the results to a central server. For instance, with a foundation model like LLaMA-3, the model can be decomposed layer-wise, enabling sequential inference by different nodes. Alternatively, decomposing the model width-wise allows for parallel inference across multiple nodes. The method of decomposition depends entirely on the application scenarios, and we consider both approaches in Nesa’s products.
2.1.3 Threat Models
@@
2.1.4 Overview of Zero-Knowledge Proofs
@@
2.1.5 Properties of Zero-Knowledge Proofs
@@ 2.1.6 Commitment
2.1.7 Proofs for Arithmetic Operations
@@
2.1.8 Proofs for Non-Arithmetic Operations
@@
2.1.9 Zero-Knowledge Proofs for Foundation Models

### v2

# Towards Secure and Private AI: A Framework for Decentralized Inference

@@ Abstract
The rapid advancement of ML models in critical sectors such as healthcare, finance, and security has intensified the need for robust data security, model integrity, and reliable outputs. Large multimodal foundational models, while crucial for complex tasks, present challenges in scalability, reliability, and potential misuse. Decentralized systems offer a solution by distributing workload and mitigating central points of failure, but they introduce risks of unauthorized access to sensitive data across nodes. We address these challenges with a comprehensive framework designed for responsible AI development. Our approach incorporates: 1) Zero-knowledge proofs for secure model verification, enhancing trust without compromising privacy. 2) Consensus-based verification checks to ensure consistent outputs across nodes, mitigating hallucinations and maintaining model integrity. 3) Split Learning techniques that segment models across different nodes, preserving data privacy by preventing full data access at any point. 4) Hardware-based security through trusted execution environments (TEEs) to protect data and computations. This framework aims to enhance security and privacy and improve the reliability and fairness of multimodal AI systems. Promoting efficient resource utilization contributes to more sustainable AI development. Our state-of-the-art proofs and principles demonstrate the framework’s effectiveness in responsibly democratizing artificial intelligence, offering a promising approach for building secure and private foundational models.
@@ 1 Introduction
Artificial Intelligence (AI), particularly machine learning (ML), has made significant strides in recent decades, with multimodal systems integrating language, vision, and audio capabilities becoming increasingly prominent. The advent of large language models (LLMs) such as ChatGPT openai2023chatgpt , Claude anthropic2023claude , Gemini google2023gemini , LLaMA llama2024 and diffusion models yang2023diffusion such as DALLE-3 openai2023dalle3 and Sora deepmind2023sora has ushered in a new era of foundation models with remarkable capabilities. While these multimodal foundation models exhibit intriguing properties such as in-context learning and chain-of-thought reasoning, they also present significant challenges in terms of security, privacy, reliability, and responsible deployment, especially in distributed or decentralized computing scenarios.
These challenges are particularly acute when considering the potential for misuse, the generation of harmful content, and the propagation of misinformation. For instance, in ML as a service (MLaaS), ensuring the integrity and reliability of inference results is paramount. Service providers must demonstrate to customers that the output stems from a verified large language model, like GPT-4, rather than from human writers or less advanced models. This requirement, referred to as model security or model integrity, is crucial for maintaining trust in AI systems. In distributed or decentralized inference applications, where a foundation model is divided into distinct segments managed by different parties, it is essential to verify the trustworthiness of each party’s execution before aggregating outputs. This distributed approach, while offering advantages in scalability and robustness, introduces new challenges in ensuring consistent and reliable outputs across nodes.
Moreover, data privacy issues are paramount in decentralized inference sun2019relationship , particularly in “critical inference” settings involving sensitive information such as medical records, financial data, or security information. For example, in healthcare applications where AI models analyze medical images for disease diagnosis, protecting patient data from unauthorized access is crucial.
We aim to address these interleaved concerns—encompassing security and privacy—for the effective deployment of AI across various domains. Our approach involves designing hybrid solutions ma2022trusted that can be tailored to specific use cases and security requirements. We discuss two key scenarios and our corresponding approaches, as well as how we adapt hardware-based trusted execution environments (TEEs) as an orthogonal approach to enhance security and privacy.
@@
Scenario 1: Critical inference refers to scenarios where the results of AI inference are extremely significant, necessitating the highest levels of security and privacy, even if it means slower speed and higher cost. In these situations, the accuracy and confidentiality of the inference outcomes are important, and users are willing to endure longer processing times to ensure their data is fully protected. Examples include healthcare diagnostics, where AI analyzes medical images to detect diseases, and financial decision-making, where AI evaluates large transactions or investment strategies. In both cases, the sensitive nature of the data demands robust privacy measures, with stakeholders prioritizing data security over rapid results to prevent unauthorized access and ensure the integrity of the process. Under this scenario, we adapt and innovate zero-knowledge machine learning (ZKML) sun2023zkdl for model integrity verification. Admittedly, the original design of ZKML is computationally heavy, especially for non-linear layers in neural networks such as ReLU layers liu2021zkcnn . To balance effectiveness and efficiency, we apply the state-of-the-art ZKML techniques introduced by sun2023zkdl ; sun2024zkllm to build zero-knowledge Decentralized Proof System (zkDPS) for proof generation and verification processes in foundation models. §2 provides more details on our customized zkDPS and SHE solutions for critical inference with the highest protection.
@@
Scenario 2: General inference refers to everyday AI inference tasks with less critical results, allowing for faster processing speeds and lower costs without compromising basic security and privacy standards. In these scenarios, a certain level of protection is still necessary, but the stringent measures required for critical inference are not needed. Examples include routine tasks such as checking the weather, recommending products, or filtering spam emails. Here, the emphasis is on efficiency and speed while maintaining adequate data security to protect user information. Under this scenario, we employ our consensus-based verification (CBV) for security, ensuring the integrity of the inference process, and split learning (SL) vepakomma2018split for data encryption, which provides a reasonable level of data protection. Notably, we have innovated this verification method to reduce the high redundancy requirements in model verification, and split learning has been used in decentralized AI encryption for the first time. §3 elaborates on our consensus-based verification and split learning solutions for general inference, balancing speed and security effectively.
@@
Our Trusted Execution Environment (TEE). In addition to our innovations around software- and/or algorithm-based security and privacy approaches, we also design an orthogonal hardware-based approach. In a nutshell, TEEs create secure isolation zones within the network’s nodes, protecting user data and private model parameters from unauthorized access sabt2015trusted . TEEs provide a secure enclave for executing computations, isolated from the rest of the node’s operating environment, ensuring that even if other parts of the node are compromised, the computations within the TEE remain protected. This isolation is critical in decentralized settings where AI models are spread among multiple owners. This hardware-based security measure works as an alternative to algorithm-based approaches, providing a fast and efficient solution for our decentralized AI inference tasks. Additionally, we innovate our TEEs by optimizing communication among multiple TEEs by establishing direct, secure channels, and implementing heterogeneous TEE scheduling based on the capabilities of each node, whether CPU or GPU-based. Our TEE facilitates secure collaboration across multiple nodes, enabling us to maintain high performance, robust security, and data privacy, making it ideal for both critical and general inference scenarios. See details of our TEE implementation and innovation in §C.
@@
Algorithm Critical (§2) zero-knowledge Decentralized Proof System (zkDPS) (§A.1) Sequential Homomorphic Encryption (SHE) (proprietary)
Algorithm General (§3) Consensus-Based Verification (CBV) (§B.1) Split Learning (SL) (§B.2)
Hardware Both TEE model verification (§C) TEE secure channel (§C)
Choosing Security and Privacy Options. Table 1 summarizes our innovations around security and privacy. Here, we further discuss the choice of different methods by scenario. Given critical inference scenarios, where the final results have high value and users are willing to wait longer, we suggest using zkDPS for model verification and SHE for data encryption. In contrast, for general inference scenarios, where the results are less critical and faster processing is desired, we recommend using CBV for model verification and SL for data encryption. Additionally, the TEE-based solution, our specialized TEE, can also be leveraged as an orthogonal approach. Once our TEE is employed, it addresses both verification and encryption needs simultaneously. However, it is important to note that GPU TEEs are only available on high-end GPUs, making the trade-off such that TEE solutions are mostly CPU-based. In contrast, algorithm-based solutions can fully leverage GPUs, offering a different balance of performance and security. This approach allows us to balance the trade-offs between security, privacy, and efficiency, ensuring that users can select the most appropriate solution for their specific needs. By leveraging both algorithm-based and hardware-based security measures, we can provide robust and adaptable security and privacy solutions across a wide range of AI applications. This adaptive approach ensures that users can select the most appropriate level of protection for their needs, achieving an optimal balance among security, privacy, and performance.
@@ 2 Security and Privacy for Critical Inference
Figure 1: Security flowchart for critical inference, where we design zkDPSA.1) for model verification and SHE (proprietary) for data encryption and user privacy protection.
@@
In critical inference scenarios, the significance of AI inference results necessitates the highest levels of security and privacy, as the outcomes directly impact crucial sectors like healthcare and finance. Our strategy for ensuring model verification and integrity is via our Zero-Knowledge Decentralized Proof System (zkDPS), a specialized ZK system for decentralized LLMs that allows one party to prove to another that a statement is true without revealing any information beyond the validity of the statement itself. Notably, it introduces a few new techniques to speed up the ZK process. This approach is detailed in §A.1.
Fig. 1 summarizes the security flow of critical inference, where the data will be transmitted with SHE for encryption, and the final inference results will be verified by zkDPS for integrity. Due to the space limitation, the extensive technical details are deferred to Appendix A.
3 Security and Privacy for General Inference
Figure 2: Security flowchart for general inference, where we design CBVB.1) for model verification and leverage SLB.2) for data encryption and user privacy protection.
In general inference scenarios, the results of AI inference tasks are less critical, allowing for faster processing speeds and lower costs without compromising basic security and privacy standards. These tasks include everyday activities such as checking the weather, recommending products, or filtering spam emails. Our strategy for ensuring model verification and integrity in these scenarios is Consensus-based Verification Check (CBV), which leverages the collective agreement of multiple nodes to ensure the correctness and integrity of model execution without revealing sensitive data. This approach is detailed in Section B.1.
For data privacy, we employ split learning (SL) vepakomma2018split ; thapa2022splitfed , a method where the model is divided into segments, and each segment is trained on different nodes to maintain data privacy by ensuring no single node has access to the complete dataset. More information can be found in Section B.2.
Fig. 2 provides an overview of the security flow of general inference in our system, which first applies SL to encrypt the first (few) layers to “encrypt” the raw data, where the AI models are sharded and deployed to nodes with redundancy. The outputs of the nodes who have the same shard will be verified by consensus for integrity. Due to the space limitation, the extensive technical details are deferred to Appendix B.
4 Conclusions and Future Directions
Artificial intelligence and machine learning have made remarkable advancements, particularly in multimodal systems integrating language, vision, and audio capabilities. While foundation models like ChatGPT, DALLE-3, and Sora demonstrate impressive capabilities, they also raise significant concerns regarding security, privacy, reliability, and responsible deployment, especially in distributed and decentralized computing scenarios. Our work addresses these challenges by offering adaptive solutions tailored to different scenarios, considering both technical and ethical implications.
We propose a comprehensive framework that ensures inference integrity and data privacy while contributing to the reliability and fairness of multimodal AI systems. For critical tasks, we recommend a zero-knowledge Decentralized Proof System (zkDPS) and Sequential Homomorphic Encryption (SHE). For general tasks prioritizing efficiency, we propose Consensus-based Verification (CBV) and Split Learning (SL). Additionally, our Trusted Execution Environment (TEE) offers a hardware-based solution addressing both verification and encryption needs. This hybrid strategy balances security, privacy, and responsible AI development, providing users with flexible, scenario-specific solutions.
Looking ahead, we aim to optimize zkDPS for real-time applications, develop adaptable ZK templates for diverse models, and create an automated framework for recommending optimal security approaches. We also plan to investigate methods to enhance the reliability of multimodal models, improve fairness, reduce bias, and develop sustainable practices for AI deployment. These efforts will focus on addressing hallucinations, misinformation propagation, and ethical considerations in decentralized systems.
By pursuing these directions, we strive to improve the robustness, efficiency, and adaptability of our solutions while contributing to the responsible development of next-generation multimodal foundational models. Our work represents a significant step towards building more secure, reliable, and ethically sound AI systems, integrating security and privacy considerations with broader responsible AI principles to foster technologies that are not only powerful and efficient but also trustworthy.
Supplementary Material
Appendix A Details of Security and Privacy for Critical Inference
In critical inference scenarios, the significance of AI inference results necessitates the highest levels of security and privacy, as the outcomes directly impact crucial sectors like healthcare and finance. Our strategy for ensuring model verification and integrity is via our Zero-Knowledge Decentralized Proof System (zkDPS), a specialized ZK system for decentralized LLMs that allows one party to prove to another that a statement is true without revealing any information beyond the validity of the statement itself. Notably, it introduces a few new techniques to speed up the ZK process. This approach is detailed in §A.1. Fig. 1 summarizes the security flow of critical inference, where the data will be transmitted with SVE for encryption, and the final inference results will be verified by zkDPS for integrity.
A.1 Zero-Knowledge Machine Learning for Model Integrity
A.1.1 Background
@@
A.1.2 Problem Setups
Decentralized Inference. We explore the challenge of decentralized inference for foundation models, which offers numerous benefits. With the advent of 5G technology and improved internet latency, personalized devices such as mobile phones can now participate in crowdsourcing machine learning models. This approach enhances device utilization and eliminates the need for data communication between nodes and a central server, thereby safeguarding data privacy. In the context of Machine Learning as a Service (MLaaS), models remain the private property of individual nodes, allowing them to offer model inference services via APIs without sharing model weights and checkpoints. As foundation models grow in size—such as Meta’s LLaMA-3.1 with 405 billion parameters [4], and future models expected to reach trillion-level sizes—it becomes impractical to load entire models on a single node or server. Additionally, technologies like blockchain are now equipped to support the decentralized inference of foundation models on-chain, further facilitating this approach.
Model Decomposition. A core assumption in decentralized AI inference is that the model used in the inference system can be divided into several parts, each managed by a separate party, or computing node. Each computing node performs its assigned computations and sends the results to a central server. For instance, with a foundation model like LLaMA-3, the model can be decomposed layer-wise, enabling sequential inference by different nodes. Alternatively, decomposing the model width-wise allows for parallel inference across multiple nodes. The method of decomposition depends entirely on the application scenarios, and we consider both approaches in our products.
A.1.3 Threat Models
@@
A.1.4 Overview of Zero-Knowledge Proofs
@@
A.1.5 Properties of Zero-Knowledge Proofs
@@ A.1.6 Commitment
A.1.7 Proofs for Arithmetic Operations
@@
A.1.8 Proofs for Non-Arithmetic Operations
@@
A.1.9 Zero-Knowledge Proofs for Foundation Models